On June 30, 2021, Georgia-Atlantic, Inc. leased a warehouse equipment from IC Leasing Corporation. The lease agreement calls for Georgia-Atlantic to make semiannual lease payments of $583,375 over a four-year lease term, payable each June 30 and December 31, with the first payment at June 30, 2021. Georgia-Atlantic's incremental borrowing rate is 12%, the same rate IC uses to calculate lease payment amounts. Amortization is recorded on a straight-line basis at the end of each fiscal year. The fair value of the equipment is $3.8 million. (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.)

Required:
1. Determine the present value of the lease payments at June 30, 2021 that Georgia-Atlantic uses to record the right-of-use asset and lease liability.
2. What pretax amounts related to the lease would Georgia-Atlantic report in its balance sheet at December 31, 2021?
3. What pretax amounts related to the lease would Georgia-Atlantic report in its income statement for the year ended December 31, 2021?
Answer #1

Solution 1:

Semiannual lease payment = $583,375

Total semiannual payments = 4*2 = 8

Incremental borrowing rate = 12%, 6% semiannual

Present value of minimum lease payments used to record right to use assets = Semi Annual lease payments * Cumulative PV Factor of annuity due for 8 periods at 6%

= $583,375 * 6.58238 = $3,839,996

Solution 2:

Semiannual payment on 30.06.2021 = $583,375

Pretax amount of liability on 30.06.2021 = ($3,839,996 - $583,375) = $3,256,621

Interest expense for 31.12.2021 = $3,256,621 * 6% = $195,397

Semiannual lease payment on 31.12.2021 = $583,375

Pre tax amount for liability December 31, 2021 = $3,256,621 + $195,397 - $583,375 = $2,868,643

Depreciation on right to use assets for 2021 = $3,839,996 / 4 * 6/12 = $480,000

Pre tax amount of right to use asset to be reported for 2021 = $3,839,996 - $480,000 = $3,359,996

Solution 3:

Pre tax amount of interest expense Georgia Atlantic Inc. reports in its income statement = $3,256,621 * 6% = $195,397

Pre tax amount of amortization expenses Georgia Atlantic Inc. reports in its income statement = $3,839,996 / 4 * 6/12 = $480,000